the sum of the digits of a three digit number is 14. The hundreds digit being 4 times the units digit. If 594 is subtracted from the number, the order of hundreds tens digits will be reversed. Find the number.
Answer by mananth(16946)

the sum of the digits of a three digit number is 14. The hundreds digit being 4 times the units digit. If 594 is subtracted from the number, the order of hundreds tens digits will be reversed. Find the number.
let the number be xyz
x+y+z=14
x=4z
Therefore 4z+y+z=14
y+5z=14..........................(1)
594 subtracted from the number
100x+10y+z-594 = 100z+10y+x
99x-99z=594
/99
x-z=6
But x=4z
4z-z=6
3z=6
z=2
if z=2 then x= 4z=>4*2=8
x+y+z=14
8+y+2=14
10+y=14
y=14-10
y=4
The number is xyz
842